OpenStudy (anonymous):

If I have an right triangle with one angle measuring 53 degrees and a side measuring in 6 inches how do I find the other two sides?

OpenStudy (anonymous):

I have already found angle 37 degrees

OpenStudy (anonymous):

Which would I use?

OpenStudy (jdoe0001):

\(\bf sin(\theta) = \cfrac{b}{c} \implies b = sin(\theta)\times c\) and so on

OpenStudy (jdoe0001):

just solve the trig identity, for the "missing side"

OpenStudy (anonymous):

So x over 6in

OpenStudy (anonymous):

Wait which one is a over b

OpenStudy (jdoe0001):

a/b ahemm, adjacent/opposite

OpenStudy (jdoe0001):

ohh, not in the picture ... yes I see hmm, well use b/a :) then